### v3

# On the Existence and Behavior of Secondary Attention Sinks

Abstract
Attention sinks are tokens, often the beginning-of-sequence (BOS) token, that receive disproportionately high attention despite limited semantic relevance. In this work, we identify a class of attention sinks, which we term secondary sinks, that differ fundamentally from the sinks studied in prior works, which we term primary sinks. While prior works have identified that tokens other than BOS can sometimes become sinks, they were found to exhibit properties analogous to the BOS token. Specifically, they emerge at the same layer, persist throughout the network and draw a large amount of attention mass. Whereas, we find the existence of secondary sinks that arise primarily in middle layers and can persist for a variable number of layers, and draw a smaller, but still significant, amount of attention mass. Through extensive experiments across 11 model families, we analyze where these secondary sinks appear, their properties, how they are formed, and their impact on the attention mechanism. Specifically, we show that: (1) these sinks are formed by specific middle-layer MLP modules; these MLPs map token representations to vectors that align with the direction of the primary sink of that layer, (2) the $\ell_{2}$-norm of these vectors determines the sink score of the secondary sink, and also the number of layers it lasts for, thereby leading to different impacts on the attention mechanisms accordingly, (3) the primary sink weakens in the middle layers, coinciding with the emergence of secondary sinks. We observe that in larger-scale models, the location and lifetime of the sinks, together referred to as sink levels, appear in a more deterministic and frequent manner. Specifically, we identify three sink levels in QwQ-32B and six levels in Qwen3-14B. We open-sourced our findings at github.com/JeffreyWong20/Secondary-Attention-Sinks.
1 Introduction
Attention sinks were first identified by Xiao et al. (2023), where the BOS token was observed to receive anomalously high attention weights. This phenomenon has since shown broad practical implications, including LLM quantization (Son et al., 2024; Liu et al., 2024), KV-cache optimization (Cai et al., 2025; 2024), efficient LLM serving (Xiao et al., 2023), and model enhancement (Yu et al., 2024).
Many recent studies investigated the formation and functional role of the BOS sink. Cancedda (2024) analyzes attention sinks from a spectral subspace perspective, while Gu et al. (2024) interprets them as a form of positional bias that mitigates over-mixing. Building on this line of work, Queipo-de-Llano et al. (2025) further examines their role in the depth-wise organization of the model, proposing that attention sinks serve as a mechanism for information compression along the depth dimension.
Sun et al. (2024) and Yu et al. (2024) show that attention sinks are not limited to the BOS token; instead, multiple tokens can function as attention sinks. Ruscio et al. (2025) further analyzes this phenomenon from a geometric perspective, showing that the emergence of multiple attention sinks is closely tied to the model’s positional embedding scheme. In this view, attention sinks act as reference points in a high-dimensional representation space, enabling the model to establish a stable internal coordinate system. However, the multiple sinks identified in previous work are fundamentally the same as the BOS sink: they emerge at the same layers and persist throughout the network. In contrast, we identify a new type of sink that differs in both its layer of emergence and its lifetime.
In this work, we show that all of the multiple tokens that act as attention sinks can be organized into distinct sink levels. The primary level normally corresponds to the BOS sink: it emerges at the same layer as the BOS token and persists throughout the network. Additional sink levels arise in the middle layers and persist for a variable number of layers; we refer to these as Secondary Sinks.
We first identify Secondary Sinks, especially their sink levels, the token sets and positions in which they frequently occur, through characterizing the similarities and differences between Secondary Sinks and the BOS sink across a range of models. (Section˜3). We then quantify the contributions of different layers to their formation via an empirical analysis of their emergence across network depth (Section˜4). Finally, we examine their impact on attention scores throughout the network.(Section˜5) . We make the following conclusions:
  • Unlike the primary sink, which emerges in early layers and persists throughout the entire network, Secondary Sinks arise primarily in middle layers and persist only for a couple of layers. They can be found at any position in the generation sequence and semantically uninformative tokens.
  • Secondary Sinks shares a similar direction with the primary sink. This direction is encoded in specific middle-layer $l_{\text{start}}$ MLP modules, which convert multiple orthogonal directions to the same sink direction. After $l_{\text{start}}$, a set of semantically uninformative tokens are transformed into attention sinks. The layers preceding $l_{\text{start}}$ play a key role in constructing this set, distinguishing these tokens from other semantically uninformative tokens.
  • Different levels of Secondary Sinks exhibit distinct lifetimes and attention sink strength. Both are strongly correlated with the $\ell_{2}$-norm of $l_{start}$ output. Larger models show clearer differentiation between sink levels, and models going through extensive post-training on reasoning data show stronger Secondary Sinks phenomenon.
  • Secondary Sinks show a compensating effect relative to the BOS sink. The BOS gradually decays and reaches its weakest strength in the middle layers, coinciding with the emergence of Secondary Sinks phenomenon.
2 Background and Preliminary
Let $f_{\theta}$ be a decoder-only transformer with $L$ layers and hidden size $h$. At each decoder layer $l$, the decoder receives a hidden sequence of length t, ${\bm{H}}^{l}\in{\mathbb{R}}^{t\times h}=\{{\bm{h}}_{0}^{l},{\bm{h}}_{1}^{l},\ldots,{\bm{h}}_{t}^{l}\}^{T}$, where ${\bm{h}}_{i}$ is the hidden representation at position $i$ in layer $l$.
Decoder blocks
Each decoder layer $l$ consists of a multi-head self-attention (MHSA) module and a multi-layer perceptron (MLP) , both of which operate on the decoder input hidden state, also referred to as the residual stream ${\bm{H}}^{l}\in{\mathbb{R}}^{t\times h}$. The MHSA produces an output ${\bm{O}}^{l}\in{\mathbb{R}}^{t\times h}$, and the MLP produces ${\bm{F}}^{l}\in{\mathbb{R}}^{t\times h}$; in both cases, the outputs are added back to the residual stream. Decoders may use either a pre-norm or post-norm architecture for which the normalization is applied either before or after each module. The majority of modern models employ pre-norm as shown in Figure˜3.
Position Embedding
Position embeddings $P$ provide tokens with positional information in the attention mechanism. Common $P$ include absolute positional embeddings (Vaswani et al., 2017), rotary positional embeddings (RoPE) (Su et al., 2024), and NTK [^1] NTK stands for Neural Tangent Kernel, this refers to changing the base frequency in RoPE based on the network capability, allowing improved performance on long-context sequences.-aware scaled RoPE (Peng et al., 2023; Xiong et al., 2024). RoPE is used in models such as LLaMA (Touvron et al., 2023) and Mistral (Jiang et al., 2023), while NTK-aware scaled RoPE is adopted by Qwen (Team and others, 2024) and CodeLlama (Roziere et al., 2023). Compared to RoPE, NTK-aware scaled RoPE increases the rotary base resulting in less rotation with position (Xiong et al., 2024).
Attention Sink
Attention sink refers to the phenomenon where semantically uninformative tokens are assigned disproportionately high attention weights across diverse model architectures and scales, resulting in vertical patterns in the attention weight matrix (Xiao et al., 2023).
Recent work Ruscio et al. (2025) suggests that reducing the rotary frequency weakens the dot-product advantage of the first token. As a result, standard RoPE often produces a strong attention sink at the BOS token, whereas NTK-aware scaled RoPE can give rise to additional attention sinks at other positions in the sequence. In this work, we evaluate 11 model families that use distinct rotary base frequencies (10K-500K). Our findings on QwQ and the Qwen2/2.5/3 model families aligns with Ruscio et al. (2025). However, we observe that CodeLlama and several other model families, despite using very large rotary bases, do not exhibit such secondary sinks. The underlying cause of this discrepancy remains an open question.
In this work, we show that while multiple tokens can act as attention sinks, they can be organised into discrete levels. The primary level corresponds to the BOS sink: these sinks emerge in the same layer as the BOS token and persist throughout the network. Additional levels appear in later layers and persist for varying depths. Building on this, we study the properties of additional sink levels, the variation of this phenomenon across different model sizes and post-training methods, then examine their formation across network depth, and finally analyze their resulting impact on attention.
3 Properties of Secondary Sinks
Figure 1: (Bottom) In LLaMA2-7B-Chat, the 1st, 11th tokens act as primary sinks: they exhibit significantly larger $l_{2}$ norms in their hidden states, while having much smaller key and value norms compared to other tokens. (Top) DeepSeek-14B also shows primary sinks (tokens 0, 1, and 2). In addition, it exhibits other attention sinks that emerge in middle layers (e.g., tokens 1120, 2768).
We perform extensive empirical experiments to study the properties of Secondary Sinks by comparing them with the BOS sink. To observe the Secondary Sinks, we generate reasoning traces from LLMs (DeepSeek-14B, DeepSeek-32B, Qwen-3-32B) on datasets including AIME24 (MAA, 2024), Math (Hendrycks et al., 2021), and then threshold the attention weights to locate sink tokens across different layers. We find that these attention sinks consistently exhibit at least an order of magnitude higher $\ell_{2}$-norms than average token [^2] Tokens excluding attention sinks. and align along the similar the BOS sink direction.
Based on this observation, we developed an efficient method to identify attention sinks by computing the pairwise cosine similarity between each hidden state ${\bm{h}}_{t!=0}^{l}$ and the BOS token ${\bm{h}}_{0}^{l}$ at layer $l$, and then thresholding with cosine similarity >> 0.95. We also observe that this phenomenon can be triggered by naturally distributed inputs rather than model-generated inputs. Accordingly, we use reasoning traces generated by DeepSeek-14B on the two datasets and feed them into 11 model families: Deepseek-distill, Qwen2, Qwen2-Math, Qwen2.5, Qwen2.5-Math, QwQ, Qwen3, LlaMA-3.1, Phi-4, Mathtral, CodeLlama. In the main body of the paper, we focus on results from DeepSeek-14B on AIME24; we observe the same behaviour in other models that exhibit secondary sinks (See Appendix Appendix˜A).
Secondary Sinks
Figure˜1 (top) shows the $\ell_{2}$-norms of the hidden states, keys, and values of attention sinks, compared with those of an average token in DeepSeek-14B denoted by $\bar{{\bm{h}}}_{t\neq\text{sink}}$. Token ${\bm{h}}_{0}$ corresponds to the BOS token, while tokens ${\bm{h}}_{1}$ and ${\bm{h}}_{2}$ are tokens from the chat template. We classify all of these sinks as primary sinks, as they emerge simultaneously and persist for the same number of layers as the BOS sink. In addition to these BOS sinks, we observe secondary sinks at positions 1120, 2768, and 9530. These tokens behave like normal tokens up to layer 22, after which they are converted into sinks. Before layer 22, their hidden-state, key, and value norms follow the average-token trend; afterward, they closely track the BOS sink, whose hidden-state norm is several orders of magnitude higher than that of normal tokens, while its key and value norms are lower.
Not all models exhibit secondary sinks. For example, although multiple sinks also appear in LLaMA-7B-Chat Figure˜1 (Bottom), they are exclusively primary sinks. Among the 11 model families we analyzed, secondary sinks were found in Qwen2, Qwen2-Math, Qwen2.5, Qwen2.5-Math, Qwen3, and QwQ. Full detection results on all 11 model families can be found in Appendix A. These sinks typically occur in semantically uninformative tokens Table˜3 and can appear at any position during the generation Figure˜5.
Figure 2: Sink levels ($l_{\text{start}}$, lifetime) at different model scales.
Sink Levels
As well as differentiating primary from secondary sinks, we can, to some extent, go further and group secondary sinks into multiple ‘levels’, based on the layer they begin in and how many layers they persist. We associate each sink with an attribute pair $(l_{\text{start}},\text{{lifetime}})$, and refer to sinks with different attribute pairs as belonging to different sink levels. We quantify the distribution of sink levels across models of different sizes and families, as shown in Figure˜2.
Secondary sinks are absent or very weak in small base models. They begin to emerge more prominently after mid-training on large amounts of mathematical data, as evidenced by the transition from Qwen2-1.5B/7B to Qwen2-Math-1.5B/7B. This suggests that secondary sinks arise as a mechanism to support enhanced reasoning capability. As model scale increases further, we observe that in the majority of models the number of distinct sink levels decreases: sink lifetimes and creation layers become less dispersed and increasingly concentrated at a small number of characteristic levels across the network.
4 Causal Formation Of Secondary Sinks
The secondary sinks exhibit an intriguing formation patterns across model layers. As shown in Figure˜1, they only become apparent as sinks after layer 22; before this layer, their hidden state $\ell_{2}$-norms are similar to those of normal tokens. To investigate their origins, we track tokens that eventually become secondary sinks, and compute their cosine similarity with the BOS sink token as they pass through the different components of layer 22. Here we focus on presenting results on Deepseek-14B on AIME24. Results on other models and datasets can be found in Appendix˜A. Our analysis reveals that these sinks primarily emerge from specific MLP modules in the network. Additionally, we perform clustering analysis on early layers and conduct token swapping experiments to trace the formation these sinks back to very early layers of the model.
Figure 3: (Left) Decoder overview (Right) Cosine Similarity Between the Different Secondary Sinks and BOS Sink Token as they go through the Layer 22 MLP in Deepseek-14B.
Cosine Similarity & PCA Analysis
For each thinking trace, we compute the cosine similarity between the hidden vector that eventually becomes a secondary sink and that of the BOS sink token at each layer. For simplicity, we name the tokens that eventually become secondary sinks as "future secondary sinks".
Figure˜3 (Right) shows the cosine similarity of 75 future secondary sink tokens from AIME24 in DeepSeek-14B (layer 22). The MLP inputs ${\bm{x}}_{t}$ initially have low similarity to the BOS direction, but this similarity increases through the MLP; by the output ${\bm{f}}_{t}$, the vectors are nearly fully aligned with the BOS sink.
This suggests that the sink direction is encoded in the MLP and that future sink tokens contain enough information to recover it. Following the linear hypothesis (Park et al., 2023), we hypothesize that these tokens share a common key that retrieves the sink direction. We test this by performing PCA on the concatenated MLP inputs ${\bm{X}}^{l}$ from 250 sink tokens. The spectrum is strongly low-rank. Feeding the top individual principal components through the MLP shows that these PCs are mapped to the same sink direction. Figure˜7 plots the MLP output norm and cosine similarity with the BOS sink when feeding $\pm\alpha\mathrm{PC}_{i}$. We find that the MLP amplifies components aligned with the BOS sink and suppresses those that are misaligned, producing the large hidden-state norms.
Clustering
After identifying the layers at which secondary sinks begin to emerge, we conduct a clustering analysis on the hidden states ${\bm{h}}_{t}^{l}$, attention outputs ${\bm{o}}_{t}^{l}$, and MLP outputs ${\bm{f}}_{t}^{l}$. These vectors correspond to two categories: (1) normal uninformative tokens and (2) future secondary sinks.
As shown in Figure˜6, in early layers, normal uninformative tokens and future sinks do not form clearly separable clusters. As the network depth increases, these two categories gradually become distinguishable, forming two increasingly distinct clusters. Both the attention and MLP modules contribute to representations that separate the two types of tokens. Notably, we observe the emergence of two well-defined clusters as early as layer 19. This shows that, although the effect of the sink doesn’t emerge until layer 22, the decision to create the sink begins in earlier layers and is mostly complete before layer 22 itself.
To better isolate the decision-making contributions from attention, MLP, and the hidden state, we substitute their outputs at future sink tokens with the average semantically uninformative token. See Appendix C for further details.
5 Secondary Sinks Effect on Attention
In this section, we first examine the difference in attention scores between secondary sinks and primary sinks. We then show that secondary sinks exhibit varying sink-score and persist for different numbers of layers. Both the sink-score and the duration across layers are correlated with the norm of the MLP output from which the sinks are formed.
Attention Sink Score
We follow Queipo-de-Llano et al. (2025) to define the attention sink-score for token position $k$ at layer $l$ as:
$\text{sink-score}_{k}^{(l,h)}=\frac{1}{T-k}\sum_{t=k}^{T-1}a_{tk}^{(l,h)},\qquad\text{sink-score}_{k}^{l}=\frac{1}{H}\sum_{h=0}^{H}\text{sink-score}_{k}^{(l,h)}$ (1)
As shown in  Figure˜4 (Left) , the sink-score for the BOS sink shows a valley shape across the model depth, and it is when it reaches its lowest sink score that the secondary sinks started emerging. This suggests the secondary sinks may serve to compensate for the decay in the BOS sink.
Figure 4: (Left) Sink-score comparison between BOS and Secondary Sinks across whole model depth. (Mid) Secondary Sinks Sink-score grows log-linearly with $\ell_{2}$-norm $l_{\text{start}}$. (Right) Sink Lifetime growths with $\ell_{2}$-norm $l_{\text{start}}$.
Attention Sink Lifetime
We observe that secondary sinks persist for varying depths, ranging from as few as 2 layers to as many as 22 layers, approximately half of the network. As shown in Figure˜4 (right), both the sink lifetime and the sink score of the secondary sink are strongly correlated with the $\ell_{2}$-norm of the MLP output ${\bm{f}}^{22}_{t}$ at $l_{\text{start}}$. The ratio between the sink scores of the secondary sink and the BOS sink exhibits a log-linear relationship with ${\bm{f}}^{22}_{t}$, with the largest slope observed at layer 33. In contrast, the sink lifetime increases monotonically with the log-norm, exhibiting distinct linear regimes separated by plateaus.
6 Conclusion
This work identifies a class of multi-token attention sinks that in certain LLMs, that differ from those described in previous works. We empirically study their properties, formation, and influence on attention patterns. We have presented evidence that secondary sinks are fundamentally different from BOS sinks. Investigating the root cause of secondary sinks emergence during pre-training, as well as why post-training processes amplify their effects, is a promising direction for future research. Moreover, its effect on the text generation and downstream performance remains an open question.
Appendix A Additional Detection Results
Model Family Model Name RoPE Base $\theta$ Primary Primary Top 3 Secondary Secondary
Sink Layer Sink Count Sink Layers Sink Count
DeepSeek deepseek-ai/DeepSeek-R1-Distill-Llama-8B 500000 1 30 X X
deepseek-ai/DeepSeek-R1-Distill-Qwen-1.5B 10000 2 60 X X
deepseek-ai/DeepSeek-R1-Distill-Qwen-7B 10000 3 60 11: 82.55% 12: 9.19% 1284
deepseek-ai/DeepSeek-R1-Distill-Qwen-14B 1000000 4 60 22: 100% 75
deepseek-ai/DeepSeek-R1-Distill-Qwen-32B 1000000 4 90 38: 98.08%, 39: 1.92% 52
Qwen3 Qwen/Qwen3-0.6B-Base 1000000 2 30 X X
Qwen/Qwen3-0.6B 1000000 2 30 X X
Qwen/Qwen3-1.7B-Base 1000000 2 30 X X
Qwen/Qwen3-1.7B 1000000 2 30 X X
Qwen/Qwen3-4B-Base 1000000 6 30 16: 100% 86
Qwen/Qwen3-4B 1000000 6 30 16: 100% 105
Qwen/Qwen3-4B-Thinking-2507 5000000 6 30 16: 100% 81
Qwen/Qwen3-4B-Instruct-2507 5000000 6 30 16: 100% 80
Qwen/Qwen3-8B-Base 1000000 6 30 16: 100% 104
Qwen/Qwen3-8B 1000000 6 30 16: 100% 134
Qwen/Qwen3-14B-Base 1000000 6 30 9: 80.51%, 8: 14.07%, 10: 3.13% 6803
Qwen/Qwen3-14B 1000000 6 30 9: 74.0%, 8: 13.85%, 38: 9.4% 7662
Qwen/Qwen3-32B 1000000 6 30 9: 38.36%, 8: 36.67%, 7: 23.06% 33727
Qwen2.5 Qwen/Qwen2.5-1.5B 1000000 2 30 X X
Qwen/Qwen2.5-1.5B-Instruct 1000000 2 30 X X
Qwen/Qwen2.5-7B 1000000 3 30 11: 50%, 12: 50% 32
Qwen/Qwen2.5-7B-Instruct 1000000 3 30 12: 56.76%, 11: 43.24% 37
Qwen/Qwen2.5-14B 1000000 5 30 22: 100% 87
Qwen/Qwen2.5-14B-Instruct 1000000 5 30 22: 100% 94
Qwen/Qwen2.5-32B 1000000 5 30 38: 98.61%, 40: 1.39% 72
Qwen/Qwen2.5-32B-Instruct 1000000 5 30 38: 100% 59
Qwen2.5-Math Qwen/Qwen2.5-Math-1.5B 10000 2 30 4: 100% 2777
Qwen/Qwen2-Math-1.5B-Instruct 10000 1 30 4: 100% 1622
Qwen/Qwen2.5-Math-7B 10000 3 30 11: 61.91%, 12: 26.34%, 13: 6.23% 95854
Qwen/Qwen2-Math-7B-Instruct 10000 3 30 12: 35.68%, 11: 29.37%, 13: 17.78% 13423
Qwen2 Qwen/Qwen2-1.5B 1000000 1 30 X X
Qwen/Qwen2-1.5B-Instruct 1000000 1 30 X X
Qwen/Qwen2-7B 1000000 3 30 17: 42.86%, 19: 14.29%, 14: 14.29% 7
Qwen/Qwen2-7B-Instruct 1000000 3 30 18: 40%, 16: 20%, 19: 20% 10
Qwen2-Math Qwen/Qwen2-Math-1.5B 10000 3 30 3: 87.64%, 4: 12.36% 356
Qwen/Qwen2-Math-1.5B-Instruct 10000 1 30 3: 100% 103
Qwen/Qwen2-Math-7B 10000 4 30 12: 33.16%, 13: 21.8%, 14: 14.94% 68709
Qwen/Qwen2-Math-7B-Instruct 10000 3 30 12: 35.53%, 13: 22.88%, 11: 15.3% 49682
Llama-3.1 meta-llama/Llama-3.1-8B 500000 1 30 X X
meta-llama/Llama-3.1-8B-Instruct 500000 1 30 X X
Phi-4 microsoft/phi-4 250000 2 30 4: 100% 30
microsoft/Phi-4-reasoning 500000 2 30 X
microsoft/phi-4-mini-reasoning 10000 3 30 X
microsoft/Phi-4-mini-instruct 10000 3 30 X
Mathstral mistralai/Mathstral-7B-v0.1 1000000 1 30 X X
Code-LLaMA meta-llama/CodeLlama-7b-hf 1000000 1 30 X X
codellama/CodeLlama-7b-Instruct-hf 1000000 1 60 X X
meta-llama/CodeLlama-13b-hf 1000000 2 30 X X
codellama/CodeLlama-13b-Instruct-hf 1000000 2 30 X X
QwQ Qwen/QwQ-32B 1000000 1 30 38: 97.37%, 40: 2.63% 38
Table 1: Model configurations and secondary sink detection results based on 30 traces generated by DeepSeek-14B on AIME24. X indicates that no secondary sinks were detected. RoPE Base $\theta$ denotes the parameter used in computing the rotary frequency; larger $\theta$ means lower frequency. Primary Sink Layer indicates the layer at which primary sinks appear. Top 3 Secondary Sink Layers denotes the top 3 creation layer $l_{\text{start}}$ at which tokens are converted into secondary sinks, with $N\%$ indicating the proportion of total sinks created by that layer. Secondary Sink Count reports the total number of detected secondary sink tokens.
Model Family Model Name RoPE Base $\theta$ Primary Primary Top 3 Secondary Secondary
Sink Layer Sink Count Sink Layers Sink Count
DeepSeek deepseek-ai/DeepSeek-R1-Distill-Llama-8B 500000 1 500 X X
deepseek-ai/DeepSeek-R1-Distill-Qwen-1.5B 10000 2 1000 X X
deepseek-ai/DeepSeek-R1-Distill-Qwen-7B 10000 3 1000 11: 69.14%, 4: 22.4%, 12: 6.39% 5807
deepseek-ai/DeepSeek-R1-Distill-Qwen-14B 1000000 4 1500 22: 99.53%, 23: 0.47% 214
deepseek-ai/DeepSeek-R1-Distill-Qwen-32B 1000000 4 1000 38: 98.55%, 39: 1.21%; 40: 0.24% 413
Qwen3 Qwen/Qwen3-0.6B-Base 1000000 2 500 X X
Qwen/Qwen3-0.6B 1000000 2 500 X X
Qwen/Qwen3-1.7B-Base 1000000 2 500 X X
Qwen/Qwen3-1.7B 1000000 2 500 X X
Qwen/Qwen3-4B-Base 1000000 6 500 16: 100% 504
Qwen/Qwen3-4B 1000000 6 500 16: 100% 692
Qwen/Qwen3-4B-Thinking-2507 5000000 6 500 16: 100% 645
Qwen/Qwen3-4B-Instruct-2507 5000000 6 500 16: 100% 623
Qwen/Qwen3-8B-Base 1000000 6 500 16: 100% 684
Qwen/Qwen3-8B 1000000 6 500 16: 100% 829
Qwen/Qwen3-14B-Base 1000000 6 500 9: 84.96%, 8: 9.37%, 10: 3.5% 36533
Qwen/Qwen3-14B 1000000 6 500 9: 81.97%, 8: 9.07%, 38: 5.56% 37349
Qwen/Qwen3-32B 1000000 6 500 8: 37.0%, 9: 34.92%, 7: 26.68% 203789
Qwen2.5 Qwen/Qwen2.5-1.5B 1000000 2 500 X X
Qwen/Qwen2.5-1.5B-Instruct 1000000 2 500 X X
Qwen/Qwen2.5-7B 1000000 3 500 11: 48.44%, 12: 48.44%, 13: 1.56% 64
Qwen/Qwen2.5-7B-Instruct 1000000 3 500 12: 54.17%, 11: 41.67%, 15: 4.17% 72
Qwen/Qwen2.5-14B 1000000 5 500 22: 99.61%, 23: 0.39% 258
Qwen/Qwen2.5-14B-Instruct 1000000 5 500 22: 100% 222
Qwen/Qwen2.5-32B 1000000 5 500 38: 98.39%, 39: 1.41%, 40: 0.2% 497
Qwen/Qwen2.5-32B-Instruct 1000000 5 500 38: 99.35%, 39: 0.65% 309
Qwen2.5-Math Qwen/Qwen2.5-Math-1.5B 10000 2 500 4: 100% 6667
Qwen/Qwen2-Math-1.5B-Instruct 10000 1 500 4: 98.74%, 1: 0.87%, 2: 0.38% 4700
Qwen/Qwen2.5-Math-7B 10000 3 500 11: 55.88%, 12: 28.87%, 13: 7.91% 149922
Qwen/Qwen2-Math-7B-Instruct 10000 3 500 12: 34.07%, 11: 29.45%, 13: 19.03% 24575
Qwen2 Qwen/Qwen2-1.5B 1000000 1 500 X X
Qwen/Qwen2-1.5B-Instruct 1000000 1 500 X X
Qwen/Qwen2-7B 1000000 3 500 17: 30.0%, 18: 30.0%, 16: 15.0% 20
Qwen/Qwen2-7B-Instruct 1000000 3 500 17: 30.0%, 18: 30.0%, 15: 15.0% 20
Qwen2-Math Qwen/Qwen2-Math-1.5B 10000 3 500 3: 93.04%, 4: 6.9%, 6: 0.06% 1739
Qwen/Qwen2-Math-1.5B-Instruct 10000 1 500 3: 99.92%, 4: 0.08% 1238
Qwen/Qwen2-Math-7B 10000 4 500 12: 30.88%, 13: 23.25%, 14: 16.26% 114579
Qwen/Qwen2-Math-7B-Instruct 10000 3 500 12: 33.96%, 13: 22.37%, 11: 17.11% 82327
Llama-3.1 meta-llama/Llama-3.1-8B 500000 1 500 X X
meta-llama/Llama-3.1-8B-Instruct 500000 1 500 X X
Phi-4 microsoft/phi-4 250000 2 500 4: 100% 500
microsoft/Phi-4-reasoning 500000 2 500 X
microsoft/phi-4-mini-reasoning 10000 3 500 X
microsoft/Phi-4-mini-instruct 10000 3 500 X
Mathstral mistralai/Mathstral-7B-v0.1 1000000 1 500 X X
Code-LLaMA meta-llama/CodeLlama-7b-hf 1000000 1 1000 X X
codellama/CodeLlama-7b-Instruct-hf 1000000 1 1000 X X
meta-llama/CodeLlama-13b-hf 1000000 2 500 X X
codellama/CodeLlama-13b-Instruct-hf 1000000 2 500 X X
QwQ Qwen/QwQ-32B 1000000 1 500 38: 95.69%, 40: 2.59%, 39: 0.86% 116
Table 2: Model configurations and secondary sink detection results based on 500 traces generated by DeepSeek-14B on AIME-500. X indicates that no secondary sinks were detected. RoPE Base $\theta$ denotes the parameter used in computing the rotary frequency; larger $\theta$ means lower frequency. Primary Sink Layer indicates the layer at which primary sinks appear. Top 3 Secondary Sink Layers denotes the top 3 creation layer $l_{\text{start}}$ at which tokens are converted into secondary sinks, with $N\%$ indicating the proportion of total sinks created by that layer. Secondary Sink Count reports the total number of detected secondary sink tokens.
Appendix B Properties of Secondary Sinks
Figure 5: Distribution of sinks across token index.
Model Secondary Sinks Token Distribution
Deepseek-14b " " "1" "^ " "2" "," "(" " I"
56.41% 17.95% 5.13% 5.13% 2.56% 2.56% 2.56%
[…]
19.87% 19.87% 13.91% 9.93% 4.64% 3.97% 3.97%
[…]
6.13% 5.28% 3.62% 3.56% 3.50% 3.48% 2.98%
Qwen2.5-14B " " "0" "1" "8" "=" "_" "t"
74.71% 9.20% 5.75% 2.30% 2.30% 1.15% 1.15%
Table 3: Frequency of secondary sink tokens.
Appendix C Causal formation of Secondary Sinks
Figure 6: Hidden State, Attention Output and MLP Output t-sne Clustering between normal semantic uninformative tokens and future sinks tokens.
Figure 7: The $\ell_{2}$-norm and cosine similarity of the MLP output relative to BOS sink with $\pm\alpha\mathrm{PC_{i}}$ as input.
Token Swapping
To further verify the formation of secondary sinks in early layers, we conduct token swapping experiments between future secondary sinks and average tokens [^3] We average the hidden vectors of all non-sink tokens at a given layer to obtain the average token representation.. By swapping the hidden vectors of these tokens, MLP outputs, and attention outputs at early layers, we observe that we can often suppress the sink from emerging in later layers. The later we swap out the activations, the more effective it is at suppressing the future sink. Figure 8 shows the percentage of the time the future sink is suppressed, as a function of the layer in which it is swapped.
Figure 8: Swapping the activation of the secondary sink at hidden state ${\bm{h}}_{t}^{l}$, attention output ${\bm{o}}_{t}^{l}$, and MLP output ${\bm{f}}_{t}^{l}$ with that of an average uninformative token.